OpenStudy (sambam):

If (-2, y) lies on the graph of y = 3x, then y =

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (sambam):

no

OpenStudy (skullpatrol):

What does the equation $$y=3x$$ mean?

OpenStudy (sambam):

I dont know?

OpenStudy (skullpatrol):

What does $$3*(-2) = -6$$ mean?

OpenStudy (skullpatrol):

or 1 + 1 = 2

OpenStudy (skullpatrol):

@sambam are you still there?

OpenStudy (sambam):

Yeah im trying to figure out what you are asking me.

OpenStudy (skullpatrol):

Is 1+1=2 true?

OpenStudy (skullpatrol):

Yes or no?

OpenStudy (sambam):

yes

OpenStudy (skullpatrol):

Is 3*(-2)=-6 true?

OpenStudy (sambam):

Oh okay yes true

OpenStudy (skullpatrol):



OpenStudy (sambam):

thanks so much.

OpenStudy (skullpatrol):

np :)